# Modern Neural Networks for Small Tabular Datasets: The New Default for Field-Scale Digital Soil Mapping?

### v1

@@ 2.2 Classical machine learning models
We included several algorithms that have demonstrated strong performance in PSM as classical ML baselines. Specifically, we evaluated Linear Regression (including regularized variants), Partial Least Squares Regression (PLSR), Random Forest, and XGBoost. Implementation details are provided in Appendix E.
Linear Regression is a statistical method that models the relationship between features and target variables using a linear equation. When multiple independent features are used, it is defined as Multiple Linear Regression (MLR). Linear Regression, specifically MLR, serves as a fundamental baseline in PSM despite its simplicity, often serving as a primary baseline for comparison against other ML algorithms (Wadoux, 2025; Ding et al., 2025; Wadoux et al., 2020). In addition to MLR, we included its regularized variants in our evaluation. Specifically, we included Lasso regression (MLR with L1 norm regularization) and Ridge regression (MLR with L2 norm regularization).
@@
Random Forest (Breiman, 2001) is an ensemble of decision trees (DT) that uses bootstrap aggregating (bagging) to reduce the variance of a single DT by employing a randomized ensemble. It is the most widely adopted ML algorithm in PSM (Ding et al., 2025; Wadoux et al., 2020). Wadoux et al. (2020) found that 80% of reviewed studies employed at least one tree-based algorithm, with Random Forest being the most popular.
XGBoost (Chen and Guestrin, 2016) is a popular implementation of GBDTs that has exhibited strong performance in many tabular data domains, including PSM. Unlike Random Forest, which trains DTs in parallel, GBDTs work by sequentially training DTs, with each new tree correcting the residuals of the previous ensemble. The optimization is performed by minimizing a differentiable loss function using gradient descent. GBDTs, and XGBoost in particular, have historically outperformed classical ML and early neural network approaches in general tabular regression benchmarks (Grinsztajn et al., 2022), although seeing more limited application in PSM compared to the widespread adoption of Random Forest (Wadoux, 2025).
@@ 2.3 Artificial neural networks
The landscape of deep learning for tabular data has evolved substantially in recent years, with various architectural innovations attempting to bridge the performance gap with traditional tree-based methods (Hollmann et al., 2025; Holzmüller et al., 2024; Gorishniy et al., 2025, 2021; Ye et al., 2025c). We included four groups of modern ANN architectures for tabular data, selecting established baseline models from each group alongside recent improvements introduced within the past year. References and details on code, implementation and dependencies of ANN models employed in this study are described in Appendix E.
@@ 2.3.1 MLP-based models
MLP represents a classical ANN baseline for tabular data. MLPs are not new to the domain of PSM and have been used and evaluated in previous studies, although generally underperforming compared to classical ML methods in field-scale PSM. We included a standard MLP as a baseline along with two recent architectural improvements, RealMLP and TabM.
@@
TabM (Gorishniy et al., 2025) extended the MLP architecture with new scaling parameter-efficient layers. The idea is to produce multiple predictions per input sample by training specialized weight-sharing components within a single model, retaining most parameters in common while still maintaining diversity through distinct prediction pathways.
@@ 2.3.4 In-context learning
We evaluated TabPFN, a foundation model that leverages in-context learning and fundamentally differs from traditional ANN approaches by eliminating the need for dataset-specific training.
TabPFN (Hollmann et al., 2025) introduced the concept of Prior-data Fitted Network (PFN) for tabular data, pre-training on a vast collection of synthetic datasets sampled from a carefully designed prior. At inference time, TabPFN processes an entire dataset, including both training and test samples, in a single forward pass, effectively performing in-context learning. This approach is claimed to be particularly effective for small to medium-sized datasets (up to 10,000 samples), where TabPFN could achieve state-of-the-art performance without any dataset-specific training. Since TabPFN was originally released as a proof-of-concept classifier without regression support (Hollmann et al., 2023), its initial application for PSM became possible only after Barkov et al. (2024) introduced a target discretization approach to adapt it to regression tasks. The most recent TabPFN version (Hollmann et al., 2025) introduced several improvements, including a native regression capability, but has not been employed or evaluated for PSM prior to the present study.
@@ 2.4.1 Fair architectural comparison
Our experimental protocol aimed to provide a fair comparison between different models. Recent advances in tabular deep learning often combine architectural innovations with complementary methodological improvements. For example, ExcelFormer introduces custom tabular-specific data augmentations, while RealMLP (Holzmüller et al., 2024) presents a comprehensive "bag-of-tricks" that extends well beyond architectural decisions to include specialized preprocessing (e.g. robust scaling with smooth clipping), postprocessing (e.g. output clipping), specific training procedures and best-epoch selection strategies. While these contributions are valuable, they are mostly model-agnostic and can theoretically be applied to any approach. We focused on unifying the evaluation protocol by either applying these improvements universally to every model, or excluding them from evaluation. This follows established conventions of fair comparison in tabular deep learning benchmarks, e.g.  Gorishniy et al. (2025) who excluded custom data augmentations when evaluating against ExcelFormer performance to maintain experimental consistency.
@@ 2.4.2 Numerical feature embeddings
PSM commonly employs continuous numerical features derived from proximal and remote sensors. Traditionally, deep learning models represent these features as raw scalars. However, Gorishniy et al. (2022) demonstrated that first transforming each scalar into a trainable vector, which they refer to as a numerical feature embedding, can significantly improve the performance of a plain MLP and allows even simple ANNs to become competitive with GBDTs on tabular benchmarks. As a consequence, modern tabular ANNs such as RealMLP, TabM, TabR and ModernNCA all implement numerical embeddings (usually piecewise-linear embeddings or their derivatives) as a core element of the network architecture. To eliminate confounding factors, we adopted a single embedding scheme for all models in this study that employ numerical feature embeddings. We used a piecewise-linear embedding module with a data-dependent number of quantile bins, implemented according to the reference code of Gorishniy et al. (2022).
@@ 2.4.3 Data preprocessing
We employed a robust scaling procedure that prevents the outliers from affecting the inlier scaling. Specifically, we preprocessed each numerical feature as follows. Let $x_{1},\dots,x_{n}\in\mathbb{R}$ be the values of a given numerical column $k$ and let $q_{p}$ be the $p$-quantile of $(x_{1},\dots,x_{n})$ for $p\in[0,1]$.
For every entry $j$ in the column $k$ we set
$\tilde{x}_{j}$ $=s_{k}\bigl{(}x_{j}-q_{1/2}\bigr{)},$
@@
For datasets whose spectroscopic feature dimensionality exceeds the number of samples we additionally applied PCA. PCA combined with MLR has repeatedly shown competitive performance in soil spectroscopy (Barra et al., 2021; Schmidinger et al., 2025). Schmidinger et al. (2025) identified PCA as the most effective preprocessing strategy when compared to using raw spectra or correlation-based feature selection in soil spectrometry. We therefore transformed the High-Dimensional datasets with PCA before model fitting, while keeping the raw variables for Low-Dimensional datasets. The number of retained components was treated as a hyperparameter and was optimized jointly with the model-specific parameters during the search described in Section 2.4.5.
@@ 2.4.4 Validation
We applied nested cross-validation with 5 outer folds for evaluation and 5 inner folds for hyperparameter selection, consistent with common ML practices in PSM (Kasraei et al., 2021).
Following previous works, (e.g. Ye et al., 2025c; Gorishniy et al., 2024), we used patience-based early stopping on the inner validation set, while keeping the inner fold test set strictly for scoring. Prior studies typically used a patience threshold of 16 consecutive epochs without validation set improvement. We extended the threshold to 40 epochs to account for the very small dataset nature in field-scale PSM, since each epoch iteration would include fewer training samples. Some architectures, such as RealMLP, train for a fixed number of epochs, typically 256, and then select the best epoch according to the validation score, which effectively mirrors a higher-patience early-stopping scheme. Following this, for final model training, we used 256 training epochs with the best epoch determined by outer validation set, while keeping the outer fold test set strictly for final evaluation. This allows us to keep the advantages of patience-based stopping serving as a run pruning strategy during hyperparameter search, and have a final model trained to escape potential early stopping patience selection bias.
@@
where $y_{i}$ are the true values, $\hat{y}_{i}$ are the predicted values, and $n$ is the number of samples.
@@ 2.4.5 Training and hyperparameter optimization
We optimized parameters of all models (including classical ML algorithms) except for TabPFN, which does not require hyperparameter tuning. For ANN architectures, we explored hyperparameters such as the number of layers, hidden dimensions, dropout rates, and additional parameters dedicated to attention or retrieval mechanisms. For tree-based methods (XGBoost, Random Forest), we tuned maximum depth, subsampling ratios, and other regularization parameters. The number of trees was set to 1,000 for Random Forest, and XGBoost employed early stopping. The detailed hyperparameter grids can be found in Appendix D.
@@ 2.4.6 Deep ensembles
We adopted deep ensembles in our evaluation framework. State-of-the-art ANN methods routinely report deep ensemble results rather than single ANN performance. For instance, FT-Transformer and TabM report results as ensembles by averaging predictions from models trained with 15 different random seeds. Similarly, ExcelFormer uses 5 runs with different seeds per dataset, while TabPFN enforces ensembling by default using 8 estimators trained with feature permutations. Beyond ensuring fair evaluation, deep ensembles have strong theoretical foundations, with recent work demonstrating that independently trained ANNs improve test accuracy when data exhibits multi-view structure (Allen-Zhu and Li, 2023) and that diversity among ensemble members is crucial for generalization (Jeffares et al., 2023b).
Furthermore, in the PSM context with extremely small datasets, we propose deep ensembles as a methodological solution to address data scarcity. Early stopping with a separate validation set requires sacrificing another portion of already limited training data. In our setup, this meant withholding 20% of the data from training set within both inner and outer folds (see Section 2.4.5). Deep ensembles enable training multiple models not only with different weight initialization and sample batching, but also on different validation splits, effectively ensuring that no single subset of data is excluded from training the deep ensemble while maintaining proper validation procedures for each individual model.
For fair evaluation across all methods, we fixed the ensemble size to 16 members for all ANN models. We conducted ablation studies to further analyze the contribution of deep ensembles to overall performance of ANNs in a PSM context (see Appendix A).
@@ 3 Results
Figure 3: Performance comparison of ANN architectures versus classical ML baselines on PSM tasks. Models are grouped by architectural approach: MLP-based (MLP, TabM, RealMLP), Attention-based (AutoInt, T2G-Former, ExcelFormer), Retrieval-based (TabR, ModernNCA), and In-Context Learning (TabPFN). Lower rank values indicate better performance. Error bars show $\pm 1$ standard error of the mean of each model’s rank across the datasets. The in-context learning model TabPFN consistently achieved the best performance across both Low-Dimensional and High-Dimensional datasets.
@@
The performance of classical ML approaches and ANN architectures assessed through average R2 ranks is shown in Figure 3. Lower rank values indicate better performance, with the best possible rank being 1 and the worst being 14 for the Low-Dimensional group and 15 for the High-Dimensional group. Since we employed different regularization schemes for MLR (see Section 2.2), the figure demonstrates the best-ranking linear model for each group. A detailed comparison of MLR, its regularized variants, and PLSR is provided in Figure B1 of Appendix B.
@@
For High-Dimensional datasets, we additionally evaluated PLSR, while all other models in the High-Dimensional group used PCA as a preprocessing step (see Section 2.4.3). The best-performing classical ML baseline for this group was MLR with Ridge penalty. While ANNs generally showed weaker absolute performance on these PCA-preprocessed datasets, most modern models still outperformed classical methods. TabPFN, similar to the Low-Dimensional group, showed the strongest performance. RealMLP ranked second after TabPFN, and retrieval-based models TabR and ModernNCA demonstrated notably stronger performance than on Low-Dimensional datasets. MLP and most attention-based models underperformed when compared to MLR with Ridge regularization.
@@
Figure 5 reveals insights about the sample size–dependent performance. To examine this, we partitioned datasets into three categories based on sample size: $\leq$50 samples, 51–120 samples, and >>120 samples. We employed two-stage ranking to avoid bias from varying numbers of models per model group. It shows that TabPFN maintained superior performance across nearly all scenarios, with classical ML surpassing it only for High-Dimensional datasets containing $\leq$50 samples. Retrieval-based models showed clear improvement with increasing dataset size in both groups, while MLP-based models exhibited stable performance regardless of size. Attention-based models showed degraded performance when PCA preprocessing was applied.
Overall, these results demonstrate that modern deep learning approaches, and TabPFN in particular, broadly surpass classical ML on the majority of field-scale PSM tasks.

### v2

@@ 2.2 Classical machine learning models
We included several algorithms that have demonstrated strong performance in pedometric modeling within DSM as classical ML baselines. Specifically, we evaluated Linear Regression (including regularized variants), Partial Least Squares Regression (PLSR), Random Forest, and XGBoost. Implementation details are provided in the Supporting Information G.
Linear Regression is a statistical method that models the relationship between features and target variables using a linear equation. When multiple independent features are used, it is defined as Multiple Linear Regression (MLR). Linear Regression, specifically MLR, serves as a fundamental baseline in DSM soil property prediction despite its simplicity, often serving as a primary baseline for comparison against other ML algorithms (Wadoux, 2025; Ding et al., 2025; Wadoux et al., 2020a). In addition to MLR, we included its regularized variants in our evaluation. Specifically, we included Lasso regression (MLR with L1 norm regularization) and Ridge regression (MLR with L2 norm regularization).
@@
Random Forest (Breiman, 2001) is an ensemble of decision trees (DT) that uses bootstrap aggregating (bagging) to reduce the variance of a single DT by employing a randomized ensemble. It is the most widely adopted ML algorithm in DSM soil property prediction (Ding et al., 2025; Wadoux et al., 2020a). Wadoux et al. (2020a) found that 80% of reviewed studies employed at least one tree-based algorithm, with Random Forest being the most popular.
XGBoost (Chen and Guestrin, 2016) is a popular implementation of GBDTs that has exhibited strong performance in many tabular data domains, including DSM. Unlike Random Forest, which trains DTs in parallel, GBDTs work by sequentially training DTs, with each new tree correcting the residuals of the previous ensemble. The optimization is performed by minimizing a differentiable loss function using gradient descent. GBDTs, and XGBoost in particular, have historically outperformed classical ML and early neural network approaches in general tabular regression benchmarks (Grinsztajn et al., 2022), although seeing more limited application in DSM compared to the widespread adoption of Random Forest (Wadoux, 2025).
@@ 2.3 Artificial neural networks
The landscape of deep learning for tabular data has evolved substantially in recent years, with various architectural innovations attempting to bridge the performance gap with traditional tree-based methods (Hollmann et al., 2025; Holzmüller et al., 2024; Gorishniy et al., 2025, 2021; Ye et al., 2025c). We included four groups of modern ANN architectures for tabular data, selecting established baseline models from each group alongside recent improvements introduced within the past year. Implementation details, model dependencies, and a concise glossary of ANN terminology are provided in the Supporting Information G and H.
@@ 2.3.1 MLP-based models
MLP represents a classical ANN baseline for tabular data. MLPs are not new to the domain of DSM and have been used and evaluated in previous studies, although generally underperforming compared to classical ML methods in field-scale DSM. We included a standard MLP as a baseline along with two recent architectural improvements, RealMLP and TabM.
@@
TabM (Gorishniy et al., 2025) extended the MLP architecture with new scaling parameter-efficient layers. The idea is to produce multiple predictions per input instance by training specialized weight-sharing components within a single model, retaining most parameters in common while still maintaining diversity through distinct prediction pathways.
@@ 2.3.4 In-context learning
We evaluated TabPFN, a foundation model that leverages in-context learning. It fundamentally differs from traditional ANN approaches by eliminating the need for dataset-specific training. Instead, it relies on pre-training on synthetic data and treats the usual training data as a context set provided to the model at inference time.
TabPFN (Hollmann et al., 2025) introduced the concept of Prior-data Fitted Network (PFN) for tabular data, pre-training on a vast collection of synthetic datasets sampled from a carefully designed prior. At inference time, TabPFN processes an entire dataset, including both training and test samples, in a single forward pass, effectively performing in-context learning. This approach is claimed to be particularly effective for small to medium-sized datasets (up to 10,000 observations), where TabPFN could achieve state-of-the-art performance without any dataset-specific training. Since TabPFN was originally released as a proof-of-concept classifier without regression support (Hollmann et al., 2023), its initial application for pedometrics and DSM became possible only after Barkov et al. (2024) introduced a target discretization approach to adapt it to regression tasks. The most recent TabPFN version (Hollmann et al., 2025) introduced several improvements, including a native regression capability, but has not been employed or evaluated for pedometric modeling within DSM prior to the present study.
@@ 2.4.1 Fair architectural comparison
Our experimental protocol aimed to provide a fair comparison between different models. Recent advances in tabular deep learning often combine architectural innovations with complementary methodological improvements. For example, ExcelFormer introduces custom tabular-specific data augmentations, while RealMLP (Holzmüller et al., 2024) presents a comprehensive "bag-of-tricks" that extends well beyond architectural decisions to include specialized preprocessing (e.g., robust scaling with smooth clipping), postprocessing (e.g., output clipping), specific training procedures and best-epoch selection strategies. While these contributions are valuable, they are mostly model-agnostic and can theoretically be applied to any approach. We focused on unifying the evaluation protocol by either applying these improvements universally to every model, or excluding them from evaluation. This follows established conventions of fair comparison in tabular deep learning benchmarks, e.g.  Gorishniy et al. (2025) who excluded custom data augmentations when evaluating against ExcelFormer performance to maintain experimental consistency.
@@ 2.4.2 Numerical feature embeddings
DSM commonly employs continuous numerical features derived from proximal and remote sensors. Traditionally, deep learning models represent these features as raw scalars. However, Gorishniy et al. (2022) demonstrated that first transforming each scalar into a trainable vector, which they refer to as a numerical feature embedding, can significantly improve the performance of a plain MLP and allows even simple ANNs to become competitive with GBDTs on tabular benchmarks. As a consequence, modern tabular ANNs such as RealMLP, TabM, TabR and ModernNCA all implement numerical embeddings (usually piecewise-linear embeddings or their derivatives) as a core element of the network architecture. To eliminate confounding factors, we adopted a single embedding scheme for all models in this study that employ numerical feature embeddings. We used a piecewise-linear embedding module with a data-dependent number of quantile bins, implemented according to the reference code of Gorishniy et al. (2022).
@@ 2.4.3 Data preprocessing
We centered and scaled each feature by subtracting the median and dividing by the interquartile range (IQR), following Holzmüller et al. (2024). This procedure is commonly referred to as robust scaling. Unlike the mean and standard deviation typically used for scaling, the median and IQR are less sensitive to outliers, yielding more stable scaling statistics.
Specifically, for each column, let $x_{1},\dots,x_{n}\in\mathbb{R}$ be the observed values and let $q_{p}$ denote the $p$-quantile of $(x_{1},\dots,x_{n})$ for $p\in[0,1]$.
For every entry $i$ we set
$\tilde{x}_{i}$ $=s\bigl(x_{i}-q_{1/2}\bigr),$
@@
For datasets whose spectroscopic feature dimensionality exceeds the number of soil samples we additionally applied PCA. PCA combined with MLR has repeatedly shown competitive performance in soil spectroscopy (Barra et al., 2021; Schmidinger et al., 2025). Schmidinger et al. (2025) identified PCA as the most effective preprocessing strategy when compared to using raw spectra or correlation-based feature selection in soil spectrometry. We therefore transformed the High-Dimensional datasets with PCA before model fitting, while keeping the raw variables for Low-Dimensional datasets. For datasets containing both spectroscopic and other sensor-derived features, PCA was applied to the combined feature set after robust scaling, thereby preserving potential covariance structures across sensor modalities. The number of retained components was treated as a hyperparameter and was optimized jointly with the model-specific parameters during the search described in Section 2.4.5.
@@ 2.4.4 Validation
We applied nested cross-validation with 5 outer folds for evaluation and 5 inner folds for hyperparameter selection, consistent with common ML practices in DSM (Kasraei et al., 2021). The complete algorithm detailing this is provided in Algorithm S1 in the Supporting Information C.
Following previous works, (e.g., Ye et al., 2025c; Gorishniy et al., 2024), we used patience-based early stopping on the inner validation set, while keeping the inner fold test set strictly for scoring. Prior studies typically used a patience threshold of 16 consecutive epochs without validation set improvement. We extended the threshold to 40 epochs to account for the very small dataset nature in field-scale DSM, since each epoch iteration would include fewer training observations. Some architectures, such as RealMLP, train for a fixed number of epochs, typically 256, and then select the best epoch according to the validation score, which effectively mirrors a higher-patience early-stopping scheme. Following this, for final model training, we used 256 training epochs with the best epoch determined by outer validation set, while keeping the outer fold test set strictly for final evaluation. This allows us to keep the advantages of patience-based stopping serving as a run pruning strategy during hyperparameter search, and have a final model trained to escape potential early stopping patience selection bias.
@@
where $y_{i}$ are the true values, $\hat{y}_{i}$ are the predicted values, and $n$ is the number of observations.
@@ 2.4.5 Training and hyperparameter optimization
We optimized parameters of all models (including classical ML algorithms) except for TabPFN, which does not require hyperparameter tuning. For ANN architectures, we explored hyperparameters such as the number of layers, hidden dimensions, dropout rates, and additional parameters dedicated to attention or retrieval mechanisms. For tree-based methods (XGBoost, Random Forest), we tuned maximum depth, subsampling ratios, and other regularization parameters. The number of trees was set to 1,000 for Random Forest, and XGBoost employed early stopping. The detailed hyperparameter grids can be found in Tables S2 and S3 in the Supporting Information F.
@@ 2.4.6 Deep ensembles
Our evaluation framework included deep ensembles to aggregate predictions from multiple independent training runs of each ANN architecture. State-of-the-art ANN methods routinely report deep ensemble results rather than single ANN performance. For instance, FT-Transformer and TabM report results as ensembles by averaging predictions from models trained with 15 different random seeds. Similarly, ExcelFormer uses 5 runs with different seeds per dataset, while TabPFN enforces ensembling by default using 8 estimators trained with feature permutations. Beyond ensuring fair evaluation, deep ensembles have strong theoretical foundations, with recent work demonstrating that independently trained ANNs improve test accuracy when data exhibits multi-view structure (Allen-Zhu and Li, 2023) and that diversity among ensemble members is crucial for generalization (Jeffares et al., 2023b).
Furthermore, in the DSM context with extremely small datasets, we propose deep ensembles as a methodological solution to address data scarcity. Early stopping with a separate validation set requires sacrificing another portion of already limited training data. In our setup, this meant withholding 20% of the data from training set within both inner and outer folds (see Section 2.4.5). Deep ensembles enable training multiple models not only with different weight initialization and data shuffling, but also on different validation splits, effectively ensuring that no single subset of data is excluded from training the deep ensemble while maintaining proper validation procedures for each individual model.
For fair evaluation across all methods, we fixed the ensemble size to 16 members for all ANN models. We conducted ablation studies to further analyze the contribution of deep ensembles to overall performance of ANNs in a DSM context (see Figure S2 in the Supporting Information B).
@@ 3 Results
Figure 3: Performance comparison of ANN architectures versus classical ML baselines on DSM tasks. Models are grouped by architectural approach: MLP-based (MLP, TabM, RealMLP), Attention-based (AutoInt, T2G-Former, ExcelFormer), Retrieval-based (TabR, ModernNCA), and In-Context Learning (TabPFN). Lower rank values indicate better performance. Error bars show $\pm 1$ standard error of the mean of each model’s rank across the datasets. The in-context learning model TabPFN consistently achieved the best performance across both Low-Dimensional and High-Dimensional datasets.
@@
The performance of classical ML approaches and ANN architectures assessed through average R2 ranks is shown in Figure 3. Lower rank values indicate better performance, with the best possible rank being 1 and the worst being 14 for the Low-Dimensional group and 15 for the High-Dimensional group. Since we employed different regularization schemes for MLR (see Section 2.2), the figure demonstrates the best-ranking linear model for each group. A detailed comparison of MLR, its regularized variants, and PLSR is provided in Figure S1 in the Supporting Information A.
@@
For High-Dimensional datasets, we additionally evaluated PLSR, while all other models in the High-Dimensional group used PCA as a preprocessing step (see Section 2.4.3). The best-performing classical ML baseline for this group was MLR with Ridge penalty. While ANNs generally showed weaker absolute performance on these PCA-preprocessed datasets, most modern models still outperformed classical methods. TabPFN, similar to the Low-Dimensional group, showed the strongest performance. RealMLP ranked second after TabPFN, and retrieval-based models TabR and ModernNCA demonstrated notably stronger performance than on Low-Dimensional datasets. MLP and most attention-based models underperformed when compared to MLR with Ridge regularization. For the distributions of absolute $R^{2}$ scores across all tasks, refer to Figure S3 in the Supporting Information D.
@@
Figure 5 reveals insights about the sample size-dependent performance. To examine this, we partitioned datasets into three categories based on sample size: $\leq$50 observations, 51–120 observations, and >>120 observations. We employed two-stage ranking to avoid bias from varying numbers of models per model group. It shows that TabPFN maintained superior performance across nearly all scenarios, with classical ML surpassing it only for High-Dimensional datasets containing $\leq$50 observations. Retrieval-based models showed clear improvement with increasing dataset size in both groups, while MLP-based models exhibited stable performance regardless of size. Attention-based models showed degraded performance when PCA preprocessing was applied.
Overall, these results demonstrate that modern deep learning approaches, and TabPFN in particular, broadly surpass classical ML on the majority of field-scale DSM tasks.